OpenStudy (anonymous):

You would like to have $1,000,000 in 33 years by making regular deposits at the end of each month in an annuity that pays 7% compounded monthly. Determine the deposit at the end of each month. Do not round until the final answer. Then round to the nearest dollar as needed.

OpenStudy (anonymous):

I know we have to use this formula FV=C((1+r)^t −1)/r But how do we use it if the payment is compounded monthly

OpenStudy (anonymous):

use FV=C((1+r/n)^t −1)/(r/n), where n is 12

OpenStudy (anonymous):

i got 648

OpenStudy (anonymous):

wow sorry gotta fix the formula i gave you FV=C((1+r/n)^(nt) −1)/(r/n) forgot a "n". sorry the answer above is correct though

OpenStudy (anonymous):

i made some kind of mistake :( so I had to retake the homework all over again so .. it doesn't really matter now because i have to do all those 10 question all over again but thank u though